# HateDebias: On the Diversity and Variability of Hate Speech Debiasing

### v1

@@ Abstract
Hate speech on social media is ubiquitous but urgently controlled. Without detecting and mitigating the biases brought by hate speech, different types of ethical problems. While a number of datasets have been proposed to address the problem of hate speech detection, these datasets seldom consider the diversity and variability of bias, making it far from real-world scenarios. To fill this gap, we propose a benchmark, named HateDebias, to analyze the model ability of hate speech detection under continuous, changing environments. Specifically, to meet the diversity of biases, we collect existing hate speech detection datasets with different types of biases. To further meet the variability (i.e., the changing of bias attributes in datasets), we reorganize datasets to follow the continuous learning setting. We evaluate the detection accuracy of models trained on the datasets with a single type of bias with the performance on the HateDebias, where a significant performance drop is observed. To provide a potential direction for debiasing, we further propose a debiasing framework based on continuous learning and bias information regularization, as well as the memory replay strategies to ensure the debiasing ability of the model. Experiment results on the proposed benchmark show that the aforementioned method can improve several baselines with a distinguished margin, highlighting its effectiveness in real-world applications.
1 Introduction
@@
Hate speech [1, 2] is frequent on social media, which has become a critical social problem and is desired to solve. One mainstream solution is to employ the technique of machine learning to detect and further mitigate hate speech [3, 4, 5], hence a safe and inclusive environment of the internet can be protected by removing these hate speeches. To well devise algorithms for hate speech detection, one of the most important steps is to provide benchmarks that are close to real-world scenarios.
@@
There are different types of benchmarks to evaluate the hate speech detection algorithm. Previously, ETHOS [6] is built from YouTube and Reddit comments validated through a crowdsourcing platform, which is designed to recognize the target of hate speech. The Measuring Hate Speech corpus [7] is devised to identify the target community and classify a sentence is a hate speech or not. In summary, these datasets are devised to recognize the target of hate speech. Recently, researchers have aimed to address the problem of debiasing and propose different datasets. Specifically, RaceBiasHate [8] uses the Twitter data annotated for hate speech and abusive language to address the racial bias. Geo-CulturalBiasHate [9] evaluates cultural bias in hate speech datasets by leveraging two interrelated cultural proxies: language and geography. Furthermore, the MultilingualBiasHate corpus [10] further considers different types of hate speech like age, country, gender, and race. In summary, existing datasets usually focus on the target of hate speech and seldom consider different types of biases.
@@
Considering the multiple types of biases of hate speech is essential to developing the hate speech debiasing models, as it prevents models from ignoring different biases during training. Although these datasets have boosted the development of hate speech detection algorithms to some extent, several algorithms trained on these datasets can hardly achieve ideal performance in practice since they are still far from real-world scenarios. As shown in Figure 1 (a), since existing datasets contain a single type of bias, the models trained on these datasets may achieve suboptimal performance on real-world scenarios with multiple and time-varying biases in Figure 1 (b). To address this challenge, it is desired to build a benchmark that close to real-world application. As shown in Figure 1 (c), when simulating how the multiple types of hate speech occur continuously, we can build a dataset with a similar distribution as the real-world scenarios, so the models may achieve ideal performance.
Table 1: Comparisons of different datasets. BI: Biased Information; D: Diversity; V: Variability.
@@
To achieve this, we first formalize the hate speech scenarios with multiple types and dynamically varying biases as continuous debiasing tasks. Our goal is to de-bias each sub-dataset with different types of hate speech biases, and meanwhile, to maintain high performance on downstream tasks. Sequentially, we construct a real-world dataset, HateDebias, to serve as a benchmark for these continuous debiasing tasks. It consists of 23276 hate speech text with 4 types of bias attributes, including age, country, gender, and ethnicity. Specifically, we simulate the emergence of various new biases at different stages based on the [11] and [10], such that each type of bias corresponds to a specific sub-dataset in HateDebias. As shown in Table 1, one can see that our dataset offers greater diversity and variability compared to existing hate speech detection datasets. Moreover, we propose a continual-learning-based framework, leveraging a bias information regularization (BIR) strategy and a memory replay (MR) strategy to better understand and tackle biases in the training process.
Contribution.
@@
(1) We explicitly formalize the continuous debiasing task that tackles the continuous debiasing scenario with multi-types and continuous-varying bias. (2) We construct a real-world dataset, called HateDebias, where any sequence of sub-dataset in HateDebias corresponds to continuous-varying bias. (3) We propose a simple yet effective framework for continuous debiasing tasks, which provides a potential direction to improve continuous debiasing tasks.
Figure 2: Overview of continuous debiasing task. The proposed HateDebias benchmark includes three aspects: task definition, dataset construction and practical framework.
2 Benchmark Continuous Debiasing Tasks
@@
In this section, we aim to benchmark the continuous debiasing task. As shown in Figure 2, the proposed HateDebias benchmark includes three aspects: task definition, dataset construction and practical framework. Specifically, we first explicitly define the continuous debiasing task that aims to continually update the model to adapt to new biases as they emerge (Sec. 2.1). Secondly, we propose a real-world dataset in which various biases are introduced periodically, used as the benchmark for evaluating continuous debiasing tasks (Sec. 2.2). Finally, we design a continual-learning-based framework incorporating techniques like bias regularization and memory replay to effectively debias over time, as a general solution for handling continuous debiasing tasks (Sec. 2.3).
2.1 Continuous Debiasing Task Definition
We focus on attribute-incremental continual learning scenarios. For example, consider a hate speech detection application used by a social media platform. Initially, the model is deployed to detect hate speech and mitigate racial bias in user-generated content. This model, while effective, must operate in a dynamic environment where new types of biases would emerge. Suppose the platform then observes an increasing trend of gender-biased hate speech. When this new bias is detected, the existing model must be adapted to address and mitigate this bias without neglecting its original capability.
@@ 2.1 Continuous Debiasing Task Definition
Let $D_{i}$ be a dataset and $A_{i}$ be the bias attribute corresponding to $D_{i}$, denoted their relationship as $BiasArr(D_{i})=A_{i}$, the continuous debias task is defined as follows:
Definition 2.1 (Continuous Debiasing Task).
@@
Given a hate speech dataset $\mathcal{D}=\{D_{1},\cdots,D_{n}\}$ with its corresponding set of bias attribute $\mathcal{A}=\{A_{1},\cdots,A_{n}\}$ that satisfies $BiasArr(D_{i})=A_{i}$, the continuous debiasing task, denote as a task sequence $\mathcal{T}_{\textit{de-bias}}=\{\operatorname{T_{1}},\cdots,\operatorname{T_{n}}\}$, aims to de-bias each sub-dataset sequentially, i.e., $\operatorname{MBias}(\operatorname{T_{i}}(D_{i},A_{i}))\neq A_{i}$, $i=1,2,\cdots,n$, where $\operatorname{MBias}(\cdot)$ is a function that reports the type of hate speech bias of a model, $\operatorname{T_{i}}$ is a sub-task that accept dataset $D_{i}$ and its corresponding bias attribute $A_{i}$ as input, and output an adjusted model.
@@
There are two key characteristics of the continuous debiasing task: diversity and variability. We will properly define them to induce the construction of the benchmark dataset. Diversity aims to emphasize the multi-type bias attributes within the dataset $\mathcal{D}$, e.g., $|\mathcal{A}|\geq 3$ where ||$\cdot$|| ⋅ | represents the dimension. On the other hand, variability represents the difference in bias attributes across various stages of the task. For instance, in a continuous debiasing task sequence $\mathcal{T}_{\textit{de-bias}}=\{\operatorname{T}_{1},\cdots,\operatorname{T}_{n}\}$, $\operatorname{T}_{i}$ aims to debias the bias attribute $A_{i}$ while $\operatorname{T}_{i+1}$ addresses the bias attribute $A_{j}$, with ${A_{i}}\neq A_{j}$.
2.2 Continuous Debiasing Dataset
@@
Diversity.
There are some hate speech datasets with various bias attributes. Huang et al. [10] gather and annotate a Twitter dataset with four types of demographic attributes, including age, country, gender, and ethnicity. Through the utilization of the Twitter streaming API [^1] https://developer.twitter.com/, they obtain tweets that were annotated with hate speech labels [12, 13], along with the corresponding user profiles. The age, race, and gender attributes of the users are inferred from analyzing their profile images through Face++ [^2] https://www.faceplusplus.com/, a computer vision API capable of estimating demographic traits. The country attribute of each user is obtained by extracting geographic information and using the Google Maps API [^3] https://maps.googleapis.com to acquire structured location data. All attributes and the hate label are converted into binary format. Thus, we select the English text dataset from the hate speech detection task corpus from the Twitter platform Huang et al. [10]. The diversity of bias attributes is guaranteed since this corpus is annotated with four types of demographic attributes, including age, country, gender, and ethnicity. In this dataset, the validation set consists of 12681 samples, and the test set comprises 12682 samples, following the dataset split by Huang et al. [10].
Table 2: Label and four attributes distribution of the training set. The row represents datasets and the column shows the attributes and labels within these datasets, and the table values detail the distribution of each attribute and label. The term “Dataset” refers to the aggregate of all four subsets.
@@
Variability.
To achieve the variability of bias types in dataset $\mathcal{D}$, our strategy is to generate sub-datasets with various types of bias and rearrange the sequence of these sub-datasets. We first collect a training set of 23,276 samples from the original training set, which consists of 59,179 samples, by removing samples with missing values, specifically those missing demographic attributes. This is because the continuous debiasing task requires the known bias attribute for its corresponding dataset. Based on the collected dataset with four bias attributes, we further partition it into four distinct subsets. To ensure distribution consistent with the collected data, the key issue in this partitioning is maintaining the same proportion of samples in each subset as in the overall dataset [14]. We address this issue using an equal proportion division method. Specifically, we randomly stratify the entire dataset based on a combination of labels (hate speech and non-hate speech) and demographic attributes (age, gender, country, and ethnicity), constraining (i) the proportion of each attribute in various sub-dataset is the same as the completed data, and (ii) there are not overlap sample between different sub-dataset. The results are shown in Table 2, there are four subsets, ‘Age’, ‘Gender’, ‘Country’, and ‘Ethnicity’. One can see that the proportion of each attribute among them is consistent with the completed data. To simulate the scenarios with continuous and changing bias types, we first assume only one explicit bias attribute in the sub-dataset and no other bias attribute can be observed. It means that, for any subset $D_{i}$, only the bias attribute $A_{i}$ is under-investigated. Furthermore, we randomly generate the sequence of these sub-datasets such that for the completed dataset $\mathcal{D}$ with its partition $\{D_{1},D_{2},D_{3},D_{4}\}$, there exists a significant changing of bias attribute.
Example.
We provide sample examples from the HateDebias Dataset in Table  3, with one sample per row. We randomly selected four examples from subsets containing 5819 samples, covering both biased and unbiased cases for each sensitive attribute. Each color sequence represents specific bias variations in the HateDebias dataset, reflecting real-world biases. The dataset includes various sensitive attributes, such as age, country, gender, and ethnicity, converted into binary format. Table  2 details the number of samples and values for each attribute. The text field contains examples from a Twitter dataset for hate speech classification, with sensitive information masked to protect privacy. The symbol ✔ indicates bias for the corresponding attribute, while ✘ indicates no bias.
Table 3: Examples of the HateDebias dataset.
@@
Age median "ah feminism is still going strong, " "at least weŕe still two young, hot, blondes." " hashtag"
@@
2.3 Continual-learning-based Framework

### v2

@@ Abstract
Hate speech frequently appears on social media platforms and urgently needs to be effectively controlled. Alleviating the bias caused by hate speech can help resolve various ethical issues. Although existing research has constructed several datasets for hate speech detection, these datasets seldom consider the diversity and variability of bias, making them far from real-world scenarios. To fill this gap, we propose a benchmark HateDebias to analyze the fairness of models under dynamically evolving environments. Specifically, to meet the diversity of biases, we collect hate speech data with different types of biases from real-world scenarios. To further simulate the variability in the real-world scenarios(i.e., the changing of bias attributes in datasets), we construct a dataset to follow the continuous learning setting and evaluate the detection accuracy of models on the HateDebias, where performance degradation indicates a significant bias toward a specific attribute. To provide a potential direction, we further propose a continual debiasing framework tailored to dynamic bias in real-world scenarios, integrating memory replay and bias information regularization to ensure the fairness of the model. Experiment results on the HateDebias benchmark reveal that our methods achieve improved performance in mitigating dynamic biases in real-world scenarios, highlighting the practicality in real-world applications.
Index Terms:
Hate Speech Debiasing, Continuous Learning, Diversity, Variability, Benchmark, Continuous Debiasing.
I Introduction
@@
Hate speech [1, 2] is frequent on social media, which has become a critical social problem and is desired to solve. One mainstream solution is to employ the technique of machine learning to detect and further mitigate hate speech [3, 4, 5], hence a safe and inclusive environment of the internet can be protected by removing these hate speeches. To advance algorithms for hate speech detection, it is important to provide benchmarks that are close to real-world scenarios.
@@
The hate speech detection task aims to identify whether a given speech is hate speech or not. There are different types of benchmarks to evaluate the hate speech detection algorithm, which can be summarized in Table I. Previously, ETHOS [6] collected data from YouTube and Reddit comments, which is designed to recognize the target of hate speech. The Measuring Hate Speech corpus [7] is devised to identify the target community and classify a sentence as hate speech or not. In summary, these datasets are devised to recognize the target of hate speech. Recently, researchers have aimed to address the problem of debiasing and propose different datasets. Specifically, RaceBiasHate [8] uses the Twitter data annotated for hate speech and abusive language to address the racial bias. Geo-CulturalBiasHate [9] evaluates cultural bias in hate speech datasets by leveraging two interrelated cultural proxies: language and geography. Furthermore, the MultilingualBiasHate corpus [10] further considers different types of hate speech, consisting of age, country, gender and race. In summary, existing datasets usually focus on the target of hate speech and seldom consider different types of biases.
@@
Considering the multiple types of biases of hate speech is essential to developing hate speech debiasing models, as it prevents models from ignoring different biases during training. Although these datasets have boosted the development of hate speech detection algorithms to some extent, several algorithms trained on these datasets can hardly achieve ideal performance in practice since they are still far from real-world scenarios. As shown in Figure 1 (a), since existing datasets contain a single type of bias, the models trained on these datasets may achieve suboptimal performance on real-world scenarios with multiple and time-varying biases in Figure 1 (b). To address this challenge, it is desired to build a benchmark that is close to a real-world application. As shown in Figure 1 (c), when simulating how the multiple types of hate speech occur continuously, we can build a dataset with a similar distribution as the real-world scenarios, so the models may achieve ideal performance.
TABLE I: Comparisons of different datasets. BI: Biased Information; D: Diversity; V: Variability.
@@
To achieve this, we first formalize the hate speech scenarios with multiple types and dynamically varying biases as continuous debiasing tasks. Our goal is to alleviate hate speech biases toward the specific attribute in each sub-dataset while maintaining promising performance on hate speech detection tasks. Sequentially, we construct a real-world dataset, HateDebias, to serve as a benchmark for these continuous debiasing tasks. It consists of 23276 hate speech texts with 4 types of bias attributes, including age, country, gender, and ethnicity. Specifically, we simulate the emergence of various new biases at different stages based on the [10] and [11], such that each type of bias corresponds to a specific sub-dataset in HateDebias. As shown in Table I, one can see that our dataset offers greater diversity and variability compared to existing hate speech detection datasets. Moreover, we propose a continual-learning-based framework, leveraging a bias information regularization (BIR) strategy and a memory replay (MR) strategy to better understand and tackle biases in the training process.
@@
Contribution. (1) We explicitly formalize the continuous debiasing task that tackles the continuous debiasing scenario with multi-types and continuous-varying bias. (2) We construct a real-world dataset, called HateDebias, where any sequence of sub-dataset in HateDebias corresponds to continuous-varying bias. (3) We propose a simple yet effective framework for continuous debiasing tasks, which provides a potential direction to improve continuous debiasing tasks.
II Related Work
Figure 2: An illustration of stastic debiasing task and continuous debiasing task. In the static debiasing task, the emergence of a new bias requires a complete re-debiasing of all existing biases, resulting in an exponential increase in both training time and costs. Conversely, in the continuous debiasing task, the emergence of a new bias requires only the fine-tuning of this specific bias, eliminating the need for re-training.
Datasets of Hata Speech Detection.
Static Debiasing.
In the field of Natural Language Processing (NLP), bias mitigation strategies fall into two primary categories. The first involves training models to reduce the influence of sensitive attributes during fine-tuning [12, 13], using techniques like projection-based [14], adversarial [15], or contrastive objectives [16]. This approach requires data specifically annotated for the sensitive attribute, with the debiasing effectiveness dependent on the external corpus quality and the potential introduction of new biases. The second category, task-agnostic training, uses general corpora to address biases, such as by identifying and removing gender biases from encoded representations [17] or enhancing model fairness through selective data exclusion [18], increased dropout rates [19], or equalizing objectives [20]. Additionally, methods like the introduction of fairness-triggering perturbations [21] and equi-tuning for group equivariance in immutable models [22] have been explored, focusing on minimizing differences between standard and debiased model features.
Continuous Debiasing.
Continual Learning.
Continual learning involves updating models with new information while retaining previous knowledge, avoiding catastrophic forgetting [25]. This process entails maintaining generic internal representations for reuse across tasks and dynamically adjusting task-specific parts. Effective strategies include memory-based methods that use prior task exemplars [26] or synthetic data [27], and regularization techniques that minimize changes in learned representations under a fixed architecture [28]. For example, Li and Hoiem [29] proposed ”learning without forgetting” (LwF), a technique that introduces regularization to align the model’s output on the current data with that of the previously trained model. Knowledge distillation [30], gradient-based [31] and expansion-based approaches [32] also mitigate forgetting, each targeting the preservation and enhancement of model adaptability over time. Furthermore, some research endeavors have shifted attention toward continual-learning-based methodologies as potential solutions to address model fairness concerns. Churamani et al. [33] introduced a pioneering application of domain-incremental learning as an effective approach for mitigating bias, specifically in the context of enhancing the fairness of facial expression recognition systems. However, there is no work focusing on the continuous debiasing task.
III Illustration of Continuous Debiasing and Static Debiasing
In our work, the term “bias” refers to the unfair predictions made by the model towards speech originating from certain groups [20, 34]. For example, when analyzing the text of a man’s speech, the model might incorrectly classify it as hate speech, influenced by the male attribute. The performance of a fair model in predicting hate speech remains consistent and unbiased, unaffected by the group attributes of the speech [35, 36]. Based on this definition of bias, the task of debiasing is traditionally categorized into two types: static debiasing and continuous debiasing.
Static Debiasing. The static debiasing is not sufficient for covering real-world scenarios. Traditionally, the static debiasing focused on adjusting specific biases, as illustrated in Fig. 2. Static debiasing requires re-debiasing all biases whenever new ones appear, consuming considerable training time and resources. Such a cost makes its application impractical in real-world environments where new biases emerge continuously. For instance, if a model is initially trained to mitigate racial bias, and then gender bias becomes apparent, the entire model needs to be retrained to address both biases simultaneously. This retraining process can be highly inefficient and resource-intensive, especially as the number and types of biases continue to grow.
Continuous Debiasing. In continuous debiasing, models in ever-changing environments are expected to adjust targetedly and dynamically to correct these continuously emerging biases, as shown at the bottom of Fig. 2. The continuous debiasing allows for incremental updates to the model to address new biases as they arise without requiring retraining from scratch. This task is significantly efficient and scalable, making it more suitable for real-world applications where the social and cultural context is continuously evolving.
Continuous debiasing plays a crucial role as it enables models to adapt to the dynamic nature of biases in the real world. However, despite its significance, there is a notable lack of benchmarks available for continuous debiasing tasks. Without proper benchmarks, it becomes difficult to gauge the effectiveness of new methods and to determine which ones are truly capable of handling the complex and evolving nature of biases in real-world scenarios. Therefore, in this paper, our goal is to propose a benchmark specifically designed for continuous debiasing tasks. We will define the continuous debiasing task and construct a comprehensive dataset that encompasses a wide range of biases and scenarios that mimic real-world variations. In addition, we develop a method that is tailored to work effectively with our proposed benchmark.
IV Benchmark Continuous Debiasing Tasks
@@
In this section, we aim to benchmark the continuous debiasing task. The proposed HateDebias benchmark includes three aspects: task definition, dataset construction and practical framework. Specifically, we first explicitly define the continuous debiasing task that aims to continually update the model to adapt to new biases as they emerge (Sec. IV-A). Secondly, we propose a real-world dataset in which various biases are introduced periodically, used as the benchmark for evaluating continuous debiasing tasks (Sec. IV-B). Finally, we design a continual-learning-based framework incorporating techniques like bias regularization and memory replay to effectively debias over time, as a general solution for handling continuous debiasing tasks (Sec. IV-C).
IV-A Continuous Debiasing Task Definition
Continuous debiasing requires that models adaptively adjust in the attribute-incremental continual learning scenarios and incrementally address newly emerging biases over time. Specifically, for hate speech detection, the model initially is deployed to detect hate speech and mitigate racial bias related to the user. In a dynamic environment, gender-biased hate speech gradually occurs on the platform. In this case, the continuous debasing task aims to develop a model that can be adapted to mitigate racial and gender biases in an evolving environment without neglecting its original capability.
@@ IV-A Continuous Debiasing Task Definition
Supposing that $D_{i}$ is a dataset and $A_{i}$ represents the bias attribute corresponding to $D_{i}$, the continuous debias task is defined as follows:
Definition IV.1 (Continuous Debiasing Task).
@@
Given a hate speech dataset $\mathcal{D}=\{D_{1},\cdots,D_{n}\}$ with its corresponding bias attribute $A_{i}\in\mathcal{A}$, where $\mathcal{A}=\{\text{Age},\text{Gender},\text{Country},\text{Ethnicity}\}$. The continuous debiasing task is a task sequence $\mathcal{T}_{\textit{debias}}=\{\operatorname{T_{1}},\cdots,\operatorname{T_{n}}\}$, aiming to debias for the attribute in each sub-dataset $D_{i}$ sequentially.
@@
There are two key characteristics of the continuous debiasing task: diversity and variability. We will properly define them to induce the construction of the benchmark dataset. Diversity aims to emphasize the multi-type bias attributes within the dataset $\mathcal{D}$, e.g., $|\mathcal{A}|\geq 3$ where $|\cdot|$ represents the dimension. On the other hand, variability represents the difference in bias attributes across various stages of the task. For instance, in a continuous debiasing task sequence $\mathcal{T}_{\textit{debias}}=\{\operatorname{T}_{1},\cdots,\operatorname{T}_{n}\}$, $\operatorname{T}_{i}$ aims to debias the bias attribute $A_{i}$ while $\operatorname{T}_{i+1}$ addresses the bias attribute $A_{j}$, with ${A_{i}}\neq A_{j}$.
IV-B Continuous Debiasing Dataset
@@
TABLE II: Label and four attributes distribution of the training set. The row represents datasets and the column shows the attributes and labels within these datasets, and the table values detail the distribution of each attribute and label. The term “Dataset” refers to the aggregate of all four subsets.
@@
TABLE III: Examples of the HateDebias dataset.
@@
Age median ”ah feminism is still going strong, ” ”at least weŕe still two young, hot, blondes.” ” hashtag”
@@
Diversity
Following the Huang et al. [10], we select the English text from the hate speech detection task corpus from the Twitter platform [10] to construct HateDebias. These texts are annotated with hate speech labels [37, 38] via the Twitter streaming API [^1] https://developer.twitter.com/, along with the corresponding user profiles by analyzing their profile images via Face++ [^2] https://www.faceplusplus.com/. We guarantee the diversity of bias attributes since this corpus contains four types of demographic attributes, including age, country, gender, and ethnicity. Specifically, the protected groups based on age are the elderly and median, the gender-based protected groups encompass males and females, the protected groups based on country involve US and non-US, and the ethnicity-based protected groups are white and non-white. In this dataset, the validation set consists of 12681 samples, and the test set comprises 12682 samples.
Variability
To achieve the variability of bias types in dataset $\mathcal{D}$, our strategy is to generate sub-datasets with various types of bias and rearrange the sequence of these sub-datasets. We first collect a training set of 23,276 samples from the original training set, which consists of 59,179 samples, by removing samples with missing values, specifically those missing demographic attributes. This is because the continuous debiasing task requires the known bias attribute for its corresponding dataset. Based on the collected dataset with four bias attributes, we further partition it into four distinct subsets. To ensure distribution consistent with the collected data, the key issue in this partitioning is maintaining the same proportion of samples in each subset as in the overall dataset [39]. We address this issue using an equal proportion division method. Specifically, we randomly stratify the entire dataset based on a combination of labels (hate speech and non-hate speech) and demographic attributes (age, gender, country, and ethnicity), constraining (i) the proportion of each attribute in various sub-dataset is the same as the completed data, and (ii) there are not overlap sample between different sub-dataset. The results are shown in Table II, there are four subsets, ‘Age’, ‘Gender’, ‘Country’, and ‘Ethnicity’. One can see that the proportion of each attribute among them is consistent with the completed data. To simulate the scenarios with continuous and changing bias types, we first assume only one explicit bias attribute in the sub-dataset and no other bias attribute can be observed. It means that, for any subset $D_{i}$, only the bias attribute $A_{i}$ is under-investigated. Furthermore, we randomly generate the sequence of these sub-datasets such that for the completed dataset $\mathcal{D}$ with its partition $\{D_{1},D_{2},D_{3},D_{4}\}$, there exists a significant changing of bias attribute.
Example
We provide sample examples from the HateDebias Dataset in Table  III, with one sample per row. We randomly selected four examples from subsets containing 5819 samples, covering both biased and unbiased cases for each sensitive attribute. Each color sequence represents specific bias variations in the HateDebias dataset, reflecting real-world biases. The dataset includes various sensitive attributes, such as age, country, gender, and ethnicity, converted into binary format. Table  II details the number of samples and values for each attribute. The text field contains examples from a Twitter dataset for hate speech classification, with sensitive information masked to protect privacy. The symbol ✔ indicates bias for the corresponding attribute, while ✘ indicates no bias.
IV-C Continual-learning-based Framework